OpenStudy (anonymous):

Find the midpoint of the line segment joining (3, -4) and (7, 2).

OpenStudy (anonymous):

do you know midpoint formula? :)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[\LARGE \left(\frac{x_1+x_2}{2},\frac{y_1+y_2}{2}\right) \]

OpenStudy (anonymous):

can you substitute

OpenStudy (anonymous):

let me try to solve it.

OpenStudy (anonymous):

5. -1?

OpenStudy (anonymous):

that's correct... well done :)

OpenStudy (anonymous):

I got this answer 3 + 7/2 , 2+ (-4)/2 10/2 , -2/2 5, -1

OpenStudy (anonymous):

that's it. \[\LARGE \begin{array}{l}\left( {\frac{{{x_1} + {x_2}}}{2},\frac{{{y_1} + {y_2}}}{2}} \right)\\\left( {\frac{{3 + 7}}{2},\frac{{ - 4 + 2}}{2}} \right)\\5, - 1\end{array}\]

OpenStudy (anonymous):

the same :)...

OpenStudy (anonymous):

Thank you but what about this formula 1/2 (3+7), 1/2 (-4 +2)? Is this also right?

OpenStudy (anonymous):

btw, I like the formula you have mentioned, it is much easier and short. Thank you so much :)

OpenStudy (anonymous):

\[\LARGE \frac12 (3+7)=\frac{3+7}{2}\] those are the same... :) you can check this out for any example : http://www.purplemath.com/modules/midpoint.htm and you're welcome :)